Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Tuesday, January 05, 1999

The following is a brief summary of the environmental monitoring activities that were completed on January 05, 1999. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

01/05/99 (Tuesday)

Segment 1 (Emidio Pump Station to Copco Road): Report for 5 January 1999: Backfill crews resumed their work near the Flying J along Peace Valley Road today. Hydro testing of the section between Flying J and the Warbler Skip was reported as successful. Work at the Warbler Skip consisted of continuing dewatering, coating, and compaction of the ROW in Lebec Road. Installation of the Tejon Ranch water pipe and PPSI fiber optic conduit moved forward today. Coating crews finished painting epoxy on welds at the California Aqueduct valves, then moved to work on wrapping the final mainline welds on Tejon Bench. Some clean up was also done along Tejon Bench. Welders worked on building the risers for the I-5/Tejon bore. Paving was done on Grapevine Road, with a compaction crew working ahead of them to prepare the ROW. Work at both Emidio and Grapevine Stations occurred today. The crew at the Grapevine Creek bore spent the day working on the ditch line, as well as removing shoring boxes on the north side. Soil was screened at the north end of the road. Welding on the 8" dewatering pipe between Emidio Pump Station and Stream Crossing #1 continued along Highway 166. Additional welding and x-ray occurred on the "leftover" pieces of pipe. Some fiber optic cable was pulled on Tejon Bench.

Segment 2B (Castaic to Sylmar) Report for 01-05-99. A crew worked to clean-up and seal the splice boxes on the Newhall Ranch property between Valencia Blvd and McBean Parkway. At Romero Canyon Road, electrical work took place in the Church St vault and valve work done at the same location. Point - to -point Electrical testing was carried out from Lake Hughes to Castaic Creek vaults. Electrical work continued at the Santa Sussana Vault. The west shoulder of the road was paved from just south of Weldon Canyon Road to the vault, where the electrical work was done.

Segment 3 (Sylmar to Los Angeles). Report for 01-05-99. Electric work took place at Balboa Ave. Tie-in activities (welding, NDT) took place just south of Balboa Ave. Electrical work occurred at Highland Ave and San Fernando Rd, and at vaults. Work at the crossing at Verdugo wash continued.

Segment 5 (Lynwood to Long Beach): Report for 1-5-99. Crews potholed for utilities and completed the bore pit at the Edison Station site near EMP 113.5. Crews working at the Tosco refinery continued fabrication and wrapping of the drain lines, progressed with installation of the station lighting and pulled the fiber optic cable into the station site from Alameda St. At Alameda St east of the Tosco refinery crews backfilled and compacted the trench for the 16'' ultra mar tie-in. Segment 4B (Lynwood to El Segundo): Crews at Lynwood station progressed with brick and electrical work. In addition, red concrete was poured over underground electrical work while crews removed the hydro-test drain lines and paved several short sections of little Alameda St. Work within the El Segundo refinery was limited to final electrical testing and clean up.

Inspection Summar

In the past 24 hours, 4 inspection event(s) were reported for 4 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page